# Cyber Attacks Detection, Prevention, and Source Localization in Digital Substation Communication using Hybrid Statistical-Deep Learning

### v1

@@ IV Cyber Attacks Detection, Prevention, and Source Localization
Fig. 1 represents the architecture of the proposed method. It can be divided into two main modules, i.e., (1) intrusion detection and prevention module, which is deployed in each IED, and (2) attack source localization module, which is deployed on a centralized location within the digital substation and receives statistical information from each IED. In the following, theoretical preliminaries and the different components and their interactions are presented.
IV-A SV Frames Arrival Time Shift
@@
As introduced in Section II, the IEC 61869-9 standard mandates strict timing specifications for SV publishers. Given that in real communication networks, the theoretical frame arrival time is affected by frame transmission, propagation, queuing, and processing delays, and time synchronization errors among SV stream publisher and subscriber devices, each frame is received at a time $F_{a}^{real}(c)\neq F_{a}^{e}(c)$, within each second. By monitoring the SV subscriber Network Interface Card (NIC), for each received frame it is possible to measure $F_{a}^{real}(c)$, and consequently, compute the difference between the theoretical and empirical frame arrival time as in (2).
$F_{as}=F_{a}^{\text{real}}(c)-F_{a}^{e}(c),\quad\forall c\in\{0,1,\dots,FS-1\}$ (2)
@@
To account for the stochasticity of the communication network infrastructure, a PDF can be estimated over $F_{as}$. Then, the modeled PDF can be used as an information advantage for intrusion detection, prevention, and attack source localization. Indeed, once that this PDF is estimated, it allows to assign a probability for each frame of being legitimate, thus allowing to discern legitimate and malicious SV frames depending on their $F_{as}$ measured at the subscriber. Regarding the choosing of the appropriate PDF to be used, it has been shown in the literature that latency in communication networks can be modeled as an Exponential distribution [26]. Unfortunately, the Exponential distribution assigns a probability equal to zero for any sample lower than the lower bound of its support and thus limits the modeling capabilities of scenarios in which a reduction of communication latency is measured due to devices time synchronization errors. To overcome these challenges, an EMG distribution is used to model $F_{as}$.
IV-B Exponentially Modified Gaussian Distribution
The EMG distribution is a probability distribution that results from the convolution of a normal distribution, $Z\sim\mathcal{N}(\mu,\sigma^{2})$, with an exponential one, $E\sim\mathrm{Exp}(\lambda)$. An EMG distribution is characterized by the following PDF:
$f(x)=\frac{\lambda}{2}e^{\frac{\lambda}{2}(2\mu+\lambda\sigma^{2}-2x)}\,\mathrm{erfc}\left(\frac{\mu+\lambda\sigma^{2}-x}{\sqrt{2}\,\sigma}\right)$ (3)
where $\mu$ and $\sigma$ are the mean and standard deviation of the normal distribution, respectively, and $\lambda={1}/\tau$ is the exponential decay of the Exponential distribution. Due to the changing operating conditions, $F_{as}$ can vary in time, and thus EMG distribution parameters need to be continuously estimated and updated using the latest received SV frames. As proven by Ali et al. [27], the Method of Moments Estimation (MME) provides the best trade-off in terms of parameters estimation accuracy and efficiency. In fact, to minimize SV frames classification latency, an accurate estimator providing a closed form solution is required. By using the MME, it is possible to equate the first three theoretical moments of the EMG distribution with the three empirical ones as in (4), (5), and (6), for the mean ($m$), variance ($s^{2}$), and skewness ($\gamma_{1}$), respectively:
@@ IV-B Exponentially Modified Gaussian Distribution
where $\hat{m}_{r}$ is the $r-th$ empirical moment computed on the observed data. As shown in Oliver et al. [28], solving for the EMG distribution parameters gives:
@@
IV-C Intrusion Detection and Prevention
To discern legitimate and malicious SV frames on a per frame basis in real-time, and to detect advanced MitM attacks, the proposed method relies on four interdependent components being executed in parallel in each IED. Moreover, two buffers, i.e., the Staging Buffer ($B_{s}$) and Accepted Frames Buffer ($B_{a}$), are used to store the latest received frames which are most likely to be legitimate and the last $k$ accepted frames, respectively.
IV-C1 Arrival PDF and Replay Protections
Once a new SV frame is received at the subscriber NIC, it is processed by the first component. First, it is checked whether frames with the same smpCnt value were already accepted in the current time period. If it is the case, the replay protection mechanism is triggered, and the frame is discarded. Then, after computing $F_{as}$ for the current frame as presented in the previous section, the EMG PDF ($\phi_{EMG}$) is used to compute the probability of the frame being legitimate ($F_{p}$), such that $F_{p}=\phi_{EMG}(F_{as})$. Subsequentially, the component checks whether in $B_{s}$ a frame with the same smpCnt value as the just received frame is present. If such a frame is already present, but the probability of it being legitimate is higher than the one computed on the last received frame, the last received frame is discarded. Otherwise, if no frame with the same smpCnt is present in $B_{s}$, or the legitimacy likelihood of the new frame is higher, the just received frame is inserted in the buffer, and the already present one is discarded. In this way, it is guaranteed that exactly $FS$ frames per second are accepted and no false positive can occur when no cyber attack is ongoing. Other than computing $F_{p}$, a frame expiration time ($F_{exp}$) is assigned to each frame added to $B_{s}$, with $F_{exp}$ computed as follows:
$F_{\text{exp}}=\begin{cases}0,&F_{as}\geq\mathbb{E}[F_{as}]^{t}\\ F_{a}+r,&F_{p}\geq\varphi_{\text{EMG}}(\mathbb{E}[F_{as}]^{t})\\ F_{a}+r+3\cdot\hat{\sigma}^{t},&\text{otherwise}\end{cases}$ (10)
@@ IV-C1 Arrival PDF and Replay Protections
The rationale behind the values assigned to $F_{exp}$ is to minimize the time spent by the frames in $B_{s}$ while guaranteeing the acceptance of the SV frames with the highest likelihood of being from the legitimate device and not from the attacker. In fact, if a frame with smpCnt equal to $c$ is received at time $F_{a}^{\prime}$, and the computed $F_{as}^{\prime}$ is equal to or greater than $\hat{\mu}^{t}+\hat{\tau}^{t}=\mathbb{E}[F_{as}]^{t}$, then, for any other frame with the same smpCnt received at a time $F_{as}^{\prime\prime}>F_{as}^{\prime}$, $\varphi_{\text{EMG}}(F_{as}^{\prime\prime})=F_{p}^{\prime\prime}$ will be lower than $\varphi_{\text{EMG}}(F_{as}^{\prime})=F_{p}^{\prime}$. Thus, the frame can be accepted immediately without any further delay. In the other cases, it cannot be guaranteed that in the future frames with higher legitimacy likelihood will not be received, and thus the received frame needs to be retained until $F_{exp}$. A graphical interpretation on the values assigned to $F_{exp}$ is provided in Fig. 2. Moreover, the pseudocode for this component is provided in Algorithm 1.
Figure 2: Graphical interpretation of $F_{exp}$ value assignment.
IV-C2 Frame Acceptance
This component is responsible for processing SV frames stored in $B_{s}$, while continuously monitoring whether the current time is higher than frames’ $F_{exp}$. If this condition holds, it implies that no additional SV frames with higher likelihood of legitimacy are expected to be received. Thus, the SV frame currently stored in $B_{s}$ can be accepted, forwarded to the protection scheme for immediate use, and added to $B_{a}$, which stores the SV frames used for updating the parameters of the EMG PDF.
@@ IV-C2 Frame Acceptance
  Inputs: SV frame $(F)$
  Variables: Variables: $B_{s}$, $F_{a}$, $F_{p}$, $F_{exp}$, frame smpCnt field value ($F_{smpCnt}$), function returning EMG PDF value at $x$ ($\varphi_{EMG}(x)$), EMG expected value ($\mathbb{E}[F_{a}^{t}]$), EMG standard deviation ($\hat{\sigma}^{t}$).
  Arrival_PDF_and_Replay_Protections($F$):
        if $|F_{as}-\mathbb{E}[F_{as}]^{t}|\geq 5\cdot\hat{\sigma}^{t}$
            discard $F$  # flooding attack mitigation
            return
        $F_{p}\leftarrow\varphi_{EMG}(F_{as})$
        if $F_{as}\geq\mathbb{E}[F_{as}]^{t}$
            $F_{exp}\leftarrow 0$
        else if $F_{p}\geq\varphi_{EMG}(\mathbb{E}[F_{as}]^{t})$
            $F_{exp}\leftarrow F_{a}+(\mathbb{E}[F_{as}]^{t}-F_{as})$
        else
            $F_{exp}\leftarrow F_{a}+\left(\mathbb{E}[F_{as}]^{t}-F_{as}\right)+3\cdot\hat{\sigma}^{t}$
        $B_{s}^{len}\leftarrow\text{length}(B_{s})$
        if $B_{s}^{len}==0$ or $B_{s}[B_{s}^{len}]_{F_{smpCnt}}<F_{smpCnt}$
            $B_{s}[B_{s}^{len}]\leftarrow[F,F_{p},F_{exp}]$
            return
        for $j\leftarrow 0,\dots,B_{s}^{len}$
            if $B_{s}[j]_{F_{smpCnt}}>F_{smpCnt}$
                $B_{s}[j]\leftarrow[F,F_{p},F_{exp}]$
                return
            if $B_{s}[j]_{F_{smpCnt}}==F_{smpCnt}$
                if $B_{s}[j]_{F_{p}}<F_{p}$
                    $B_{s}[j]\leftarrow[F,F_{p},F_{exp}]$
                else
                    discard $F$
                    return
IV-C3 EMG Handler and PDF Estimator
After the number of SV frames in $B_{a}$ is greater than $k$, this component takes care of estimating and updating the parameters of the EMG distribution. This is done by first updating the samples’ moments with a weighted mean between the previously calculated moments ($\hat{m}^{t}_{r}$) and the just computed ones ($\underline{\hat{m}_{r}}$), as in (11), and then using the updated moments values to compute the updated EMG parameters $\hat{\mu}^{(t+1)}$, $\hat{\sigma}^{(t+1)}$, and $\hat{\lambda}^{(t+1)}$.
$\hat{m}_{r}^{(t+1)}=\frac{FS\cdot\hat{m}_{r}^{t}+k\cdot\underline{\hat{m}}_{r}}{FS+k}$ (11)
@@ IV-C4 MitM Detection
Whereas the first component is meant to accept exactly $FS$ SV frames per second while reducing the likelihood of malicious frames being accepted, this third component aims at detecting advanced MitM attacks. In this type of attack, it is assumed that the attacker can completely block the legitimate SV stream and inject a malicious one. Thus, the subscriber device receives exactly $FS$ frames per second as expected. Still, the act of intercepting and tampering with the forwarded SV frames causes an inevitable change in the statistics of the received SV stream at the subscriber level. In fact, to perform the MitM attack, the attacker is required to change the NIC configuration of the compromised forwarding device to disable the HSR bridge between the two NICs. Then, the network traffic needs to be monitored, tampered with, and forwarded from one NIC to the other at the device application layer, causing the inevitable alteration of the forwarded traffic statistical properties. To detect these changes, a Recurrent Neural Network (RNN) is deployed to detect anomalous variations in the EMG parameters. The input of the RNN consists of the mean ($m$), standard deviation ($s$), and skewness ($\gamma_{1}$) of the $F_{as}$ measured on the accepted frames, calculated over a sliding window of 200 samples with a step size of 50 samples. The RNN outputs the likelihood of an ongoing MitM attack. A grid search approach is used to select the most appropriate RNN model and perform hyperparameter optimization. The considered RNN models include Elman RNN, Long-Short Term Memory (LSTM), and Gated Recurrent Units (GRU), whereas the dimension and the number of hidden layers vary from 3 to 20 and from 2 to 8, respectively. The model providing the best trade-off between model complexity, detection accuracy, and prediction time consists of 4 stacked LSTM cells of 20 neurons each. Finally, to perform the binary classification, a fully connected layer is stacked to the output of the last LSTM cell.
IV-D Attack Source Localization
@@
As mentioned in the previous section, anomalous changes in the PDF of $F_{as}$ can be an indicator of an ongoing MitM attack. By aggregating and correlating the statistical properties of the SV frames arrival times from different IEDs deployed within the digital substation it is possible to identify and locate which device is performing a MitM attack. In this paper, the focus is on IEDs deployed in a HSR ring topology. As represented in Fig. 1, the mean ($m_{i}$), standard deviation ($s_{i}$), and skewness ($\gamma_{(1,i)}$) of the measured $F_{as}$ from each $i-th$ IED are delivered to the server hosting the attack source localization module within the digital substation. These metrics are then used to estimate the probability of a MitM attack being ongoing, and the probability for each IED of acting maliciously. As for the MitM component, a sliding window of 200 $F_{as}$ samples with a step size of 50 samples is used. This last component consists of 5 stacked Elman RNN cells; each cell hidden layer contains 26 neurons. The tanh activation function is used. The output of the last RNN cell is provided as input to a fully connected layer to perform the final multi-class classification. To improve the DL model accuracy and increase its generalization capabilities, the component’s input data is derived from the statistical metrics received from each IED as follows:
$\forall\,t\in\{1,\dots,T\},\,i,j\in\{1,\dots,N\},\,i<j:$
$\begin{cases}m_{(i|j)}^{t}=m_{i}^{t}-m_{j}^{t}\\ s_{(i|j)}^{t}=s_{i}^{t}-s_{j}^{t}\\ \gamma_{(1,i|j)}^{t}=\gamma_{(1,i)}^{t}-\gamma_{(1,j)}^{t}\end{cases}$ (12)
@@
where $t$ is the current time window, $N$ is the number of IEDs deployed in the HSR ring, and $m_{n}^{t}$, $s_{n}^{t}$, and $\gamma_{(1,n)}^{t}$ are the mean, standard deviation, and skewness measured at $n-th$ IED at time $t$. This data transformation prevents the DL model from overfitting the absolute values measured in the specific communication network conditions, but to focus more on the relative evolution of the statistical properties.
IV-E Attacker Considerations

### v2

@@ III Cyber Attacks Detection, Prevention, and Source Localization
The proposed integrated method can be divided into two main modules, i.e., (1) the intrusion detection and prevention module, which is deployed in each IED, and (2) the attack source localization module, which is deployed on a centralized location within the digital substation and receives statistical information from each IED. Fig. 1 represents the architecture of the proposed method. In the following, theoretical preliminaries and the different components and their interactions are presented.
III-A SV Frames Arrival Time Shift
@@
As introduced in Section II, the IEC 61869-9 standard mandates strict timing specifications for SV publishers. Given that in real communication networks, the theoretical frame arrival time is affected by frame transmission, propagation, queuing, and processing delays, and time synchronization errors among SV stream publisher and subscriber devices, each frame is received at a time $F_{a}^{r}(c)\neq F_{a}^{e}(c)$, within each second. By monitoring the SV subscriber’s Network Interface Card (NIC), the arrival time $F_{a}^{r}(c)$ of each received frame can be measured, and consequently, the difference between the theoretical and empirical frame arrival times can be computed as shown in (2).
$F_{as}=F_{a}^{r}(c)-F_{a}^{e}(c),\quad\forall c\in\{0,1,\dots,FS-1\}$ (2)
To account for the stochasticity of the communication network infrastructure, a PDF can be estimated over $F_{as}$. Then, the modeled PDF can be used as an information advantage for intrusion detection, prevention, and attack source localization. Indeed, once this PDF is estimated, a legitimacy probability can be assigned to each received frame, allowing the distinction between legitimate and malicious SV frames based on their $F_{as}$ measured at the subscriber. Regarding the choice of the most appropriate PDF to be used, the literature shows that latency in communication networks is often modeled using an exponential distribution [2]. However, the exponential distribution assigns a probability equal to zero to samples below its lower bound, making it unsuitable for scenarios where latency reductions may occur due to device time-synchronization errors. To address this limitation, the EMG distribution is used to model $F_{as}$. The EMG, defined as the sum of a Gaussian and an exponential component, effectively represents both the narrow, near-deterministic base communication delay and the heavy-tailed positive jitter caused by queueing, contention, and background traffic. Such characteristics enable the EMG to capture the sharp mode and long right tail observed in empirical network latency distributions more accurately than either distribution alone.
Finally, it should be remarked that, while related, the inter-frame arrival time, i.e., I-FAT, and SV frames arrival time shift, i.e., $F_{as}$, differ fundamentally in definition, probability distribution, information content, and susceptibility to communication network latency and jitter. In fact, I-FAT represents the time difference between consecutive SV frames and follows a Gaussian distribution centered at $1/FS$, whereas $F_{as}$ measures the deviation between the expected and actual arrival times for individual frames, modeled by an EMG distribution whose mean reflects communication network latency and clock synchronization drift [17]. These distinctions represent a fundamental difference compared to how the SV frame arrival times information has been exploited in existing studies.
@@
III-B Exponentially Modified Gaussian Distribution
The EMG distribution is a probability distribution whose PDF is obtained by convoluting a normal distribution, $Z\sim\mathcal{N}(\mu,\sigma^{2})$, with an exponential one, $E\sim\mathrm{Exp}(\lambda)$. An EMG distribution is characterized by the following PDF:
$f_{EMG}(x)=\frac{\lambda}{2}e^{\frac{\lambda}{2}(2\mu+\lambda\sigma^{2}-2x)}\,\mathrm{erfc}\left(\frac{\mu+\lambda\sigma^{2}-x}{\sqrt{2}\,\sigma}\right)$ (3)
where $\mu$ and $\sigma$ are the mean and standard deviation of the normal distribution, respectively, and $\lambda={1}/\tau$ is the exponential decay of the exponential distribution. Due to the changing operating conditions, $F_{as}$ can vary in time, and thus EMG distribution parameters need to be continuously estimated and updated using the latest received SV frames. As shown by Ali et al. [1], the Method of Moments Estimation (MME) provides the best trade-off in terms of parameters estimation accuracy and efficiency. In fact, to minimize SV frames classification latency, an accurate estimator providing a closed form solution is required. By using the MME, it is possible to equate the first three theoretical moments of the EMG distribution with the three empirical ones as in (4), (5), and (6), for the mean ($m$), variance ($s^{2}$), and skewness ($\gamma_{1}$), respectively:
@@ III-B Exponentially Modified Gaussian Distribution
where $\hat{m}_{r}$ is the rthr-th empirical moment computed on the observed data. As shown in Oliver et al. [21], solving for the EMG distribution parameters gives:
@@
III-C Intrusion Detection and Prevention
To discern legitimate and malicious SV frames on a per-frame basis in real-time, and to detect MitM attacks, the proposed method relies on four interdependent components being executed in parallel in each IED. Moreover, two buffers, i.e., the staging buffer ($B_{s}$) and the accepted frames buffer ($B_{a}$), are used to store the latest received frames which are most likely to be legitimate, and the last $k$ accepted frames, respectively.
Figure 2: Graphical interpretation of $F_{exp}$ value assignment.
III-C1 Arrival PDF and Replay Protections
Once a new SV frame is received at the subscriber NIC, it is processed by the first component. First, it is checked whether frames with the same smpCnt value were already accepted in the current time period. If it is the case, the replay protection mechanism is triggered, and the frame is discarded. Then, after computing $F_{as}$ for the current frame as presented in the previous section, the EMG PDF ($f_{EMG}$) is used to compute the probability of the frame being legitimate ($F_{p}$), such that $F_{p}=f_{EMG}(F_{as})$. Subsequently, the component checks whether in $B_{s}$ a frame with the same smpCnt value as the just received frame is present. If such a frame is already present, but the probability of it being legitimate is higher than the one computed on the last received frame, the last received frame is discarded. Otherwise, if no frame with the same smpCnt is present in $B_{s}$, or the legitimacy likelihood of the new frame is higher, the just received frame is inserted in the buffer, and the already present one is discarded. In this way, it is ensured that exactly $FS$ frames per second are accepted and no false positives can occur when no cyber attack is ongoing. Other than computing $F_{p}$, a frame expiration time ($F_{exp}$) is assigned to each frame added to $B_{s}$, with $F_{exp}$ computed as follows:
$F_{exp}=\begin{cases}0,&F_{as}\geq\mathbb{E}[F_{as}]^{t}\\ F_{a}+r,&F_{p}\geq f_{EMG}(\mathbb{E}[F_{as}]^{t})\\ F_{a}+r+3\cdot\hat{\sigma}^{t},&\text{otherwise}\end{cases}$ (10)
@@ III-C1 Arrival PDF and Replay Protections
The rationale behind the values assigned to $F_{exp}$ is to minimize the time frames remain buffered in $B_{s}$ while ensuring that frames with the highest likelihood of originating from the legitimate device, rather than from the attacker, are accepted. In fact, if a frame with smpCnt equal to $c$ is received at time $F_{a}^{\prime}$, and the computed $F_{as}^{\prime}$ is equal to or greater than $\hat{\mu}^{t}+\hat{\tau}^{t}=\mathbb{E}[F_{as}]^{t}$, then, for any other frame with the same smpCnt received at a time $F_{as}^{\prime\prime}>F_{as}^{\prime}$, $f_{EMG}(F_{as}^{\prime\prime})=F_{p}^{\prime\prime}$ will be lower than $f_{EMG}(F_{as}^{\prime})=F_{p}^{\prime}$. Thus, the frame can be accepted immediately without any further delay. In the other cases, it cannot be inferred that frames with higher legitimacy likelihood will not be received in the future, and thus the received frame needs to be retained until $F_{exp}$. A graphical interpretation on the values assigned to $F_{exp}$ is provided in Fig. 2. Moreover, the pseudocode for this component is provided in Algorithm 1.
III-C2 Frame Acceptance
This component is responsible for processing SV frames stored in $B_{s}$, while continuously monitoring whether the current time is higher than frames’ $F_{exp}$. If this condition holds, it implies that no additional SV frames with the same smpCnt field value and a higher likelihood of legitimacy are expected to be received. Thus, the SV frame currently stored in $B_{s}$ can be accepted, forwarded to the protection scheme for immediate use, and added to $B_{a}$, which stores the SV frames used for updating the parameters of the EMG PDF.
1
2
Input : SV frame $F$, Frame arrival time shift ($F_{as}$), Current buffer $B_{s}$
Output : Updated buffer $B_{s}$ or Discard frame
3
4if $|F_{as}-\mathbb{E}[F_{as}]^{t}|\geq 3\cdot\hat{\sigma}^{t}$ then
5 Discard $F$
6 Return
7
[…]
9
10if $F_{as}\geq\mathbb{E}[F_{as}]^{t}$ then
11 $F_{exp}\leftarrow 0$
12 else if $F_{p}\geq f_{\text{EMG}}(\mathbb{E}[F_{as}]^{t})$ then
13 $F_{exp}\leftarrow F_{a}+(\mathbb{E}[F_{as}]^{t}-F_{as})$
14
15 else
16 $F_{exp}\leftarrow F_{a}+(\mathbb{E}[F_{as}]^{t}-F_{as})+3\cdot\hat{\sigma}^{t}$
17
18
[…]
21 if $F^{\prime}.F_{p}<F_{p}$ then
22      Replace $F^{\prime}$ with $F_{new}$ in $B_{s}$
23
24 else
25    Discard $F$
26
27else
28   Insert $F_{new}$ into $B_{s}$ maintaining order by smpCntsmpCnt
29
30
Return
@@ 28   Insert $F_{new}$ into $B_{s}$ maintaining order by smpCntsmpCnt
III-C3 EMG Handler and PDF Estimator
After the number of SV frames in $B_{a}$ is greater than $k$, this component takes care of estimating and updating the parameters of the EMG distribution. This is done by first updating the samples’ moments with a weighted mean between the previously calculated moments ($\hat{m}^{t}_{r}$) and the just computed ones ($\underline{\hat{m}_{r}}$), as in (11), and then using the updated moments values to compute the updated EMG parameters $\hat{\mu}^{t+1}$, $\hat{\sigma}^{t+1}$, and $\hat{\lambda}^{t+1}$.
$\hat{m}_{r}^{t+1}=\frac{FS\cdot\hat{m}_{r}^{t}+k\cdot\underline{\hat{m}}_{r}}{FS+k}$ (11)
@@ III-C4 MitM Detection
Whereas the first component is meant to accept exactly $FS$ SV frames per second while minimizing the likelihood of malicious frames being accepted, this fourth component aims at detecting MitM attacks on SV streams. In this type of attack, the attacker can completely block the legitimate SV stream and inject a malicious one by compromising one forwarding device between the SV publisher and subscriber. Thus, the subscriber device receives exactly $FS$ frames per second as expected. Still, as experimentally shown, the act of intercepting and tampering with the forwarded SV frames causes an inevitable change in the latency and jitter statistics of the received SV stream at the subscriber level. In fact, to perform the MitM attack, the attacker is required to change the NIC configuration of the compromised forwarding device to modify the frames forwarding rules or to disable the HSR bridge between the two NICs. Then, the network traffic needs to be monitored, tampered with, and forwarded from one NIC to the other in user-space level on the device, causing the inevitable alteration of the forwarded traffic statistical properties. A RNN is deployed to detect anomalous variations in the EMG parameters. The input of the RNN consists of the mean ($m$), standard deviation ($s$), and skewness ($\gamma_{1}$) of the $F_{as}$ measured on the accepted frames, calculated over a sliding window of 200 samples with a step size of 50 samples. The RNN outputs the likelihood of an ongoing MitM attack. A grid search approach is used to select the most appropriate RNN model and perform hyperparameter optimization. The considered RNN models include Elman RNN, LSTM, and Gated Recurrent Units (GRU), whereas the dimension and the number of hidden layers vary from 3 to 20 and from 2 to 8, respectively. The model providing the best trade-off between model complexity, detection accuracy, and prediction time consists of 4 stacked LSTM cells of 20 neurons each. Finally, to perform the binary classification, a fully connected layer is stacked to the output of the last LSTM cell.
III-D Attack Source Localization
@@
As mentioned in the previous section, anomalous changes in the PDF of $F_{as}$ can be an indicator of an ongoing MitM attack. By aggregating and correlating the statistical properties of the SV frames arrival times from different IEDs deployed within the digital substation, it is possible to identify and locate which device is performing a MitM attack. In this paper, the focus is on IEDs deployed in an HSR ring topology. As represented in Fig. 1, the mean ($m_{i}$), standard deviation ($s_{i}$), and skewness ($\gamma_{1,i}$) of the measured $F_{as}$ from each ithi-th IED are delivered to the server hosting the attack source localization module within the digital substation. These metrics are then used to estimate the probability of a MitM attack being ongoing, and the probability for each IED of acting maliciously. As for the MitM component, a sliding window of 200 $F_{as}$ samples with a step size of 50 samples is used. This last component consists of 5 stacked Elman RNN cells; each cell hidden layer contains 26 neurons. The tanh activation function is used. The output of the last RNN cell is provided as input to a fully connected layer to perform the final multi-class classification. To improve the DL model accuracy and increase its generalization capabilities, the component’s input data is derived from the statistical metrics received from each IED as follows:
$\forall t\in\{1,\dots,T\},\forall i,j\in\{1,\dots,N\},\,i<j:$
$\begin{cases}m_{i,j}^{t}=m_{i}^{t}-m_{j}^{t}\\ s_{i,j}^{t}=s_{i}^{t}-s_{j}^{t}\\ \gamma_{1,i,j}^{t}=\gamma_{1,i}^{t}-\gamma_{1,j}^{t}\end{cases}$ (12)
@@
where $t$ is the current time window, $N$ is the number of IEDs deployed in the HSR ring, and $m_{n}^{t}$, $s_{n}^{t}$, and $\gamma_{1,n}^{t}$ are the mean, standard deviation, and skewness measured at nthn-th IED at time $t$. This data transformation prevents the DL model from overfitting the absolute values measured in the specific communication network conditions, but to focus on the relative evolution of the statistical properties.
III-E Attacker Considerations